OpenStudy (anonymous):

Medal for whoever helps. 1.Match each of the addition expressions with its rule. A.Subtract and the sum is positive. B.Add and the sum is positive. C.Add and the sum is negative. D.Subtract and the sum is negative. *3/8 + 1/3 *-3/8 + -1/2 *-5/6 + 1/2 *14/2 + -1/2 2.If a rancher sold 1/3 of his cattle and then sold another 1/4 of the original amount, what part of his cattle was sold? A.1/7 B.2/12 C.7/12 3.Harriet had a 2 1/4-foot length of board. She will use 1 2/3 feet for a project today. How many feet of board will remain? (Hint: you can use 9/4 and 5/3 if you wish.) A.7/12 B.2/7 C.14/7

jhonyy9 (jhonyy9):

1. 3/8 +1/3 =(9+8)/24=17/24 2. -3/8 + -1/2 =(-3-4)/8=-7/8 3. -5/6 +1/2 =(-5+3)/6=(-2)/6=-1/3 4. 14/2+ - 1/2 =13/2

jhonyy9 (jhonyy9):

2. 1/3 + 1/4 =7/12 -- C.

jhonyy9 (jhonyy9):

9/4 - 5/3 = (27-20)/12=7/12 --- A.